# LoL-PIM: Long-Context LLM Decoding with Scalable DRAM-PIM System

### v1

@@ 8.1. Evaluation Settings
Table 6. Validation of our Ramulator-based PIM simulator (cycle difference with the AiMX vendor library).
@@
Table 7. Specification of GPU-HBM, GPU-GDDR, and PIM(=LoL-PIM).
@@
Baselines. We evaluate our system against three baselines: (1) a 16-GPU DGX-A100 system, each GPU equipped with 80GB HBM3 and simulated using a DGX simulator (GPU-HBM), (2) 16 AiMX cards, each comprising 8 devices with one Xilinx Virtex UltraScale+ VU9P FPGA and eight 1GB AiM chips utilizing GDDR6 and QSFP links for inter-card communication (PIM), and (3) a GPU-only system using GDDR6 for GPU evaluation with the matched external bandwidth of the PIM system (GPU-GDDR). These baselines were simulated using validated tools, including the DGX simulator (Jaewan Choi, 2024) and our extended version of Ramulator (Kim et al., 2016), which we validated to ensure accuracy (see Table 6). This setup enables a comprehensive comparison of LLM inference performance across architectures. Detailed specifications for the baselines are provided in Table 7.
@@ 8.2. Performance Comparison
Scalability. We evaluate LoL-PIM’s scalability across three long-context LLM benchmarks(Table 2), focusing on the Qwen1.5-7B model (Fig. 9(a)). The baseline PIM-only system showed the poorest scalability due to inefficient PIM utilization. While it outperformed the GPU-only system at a smaller capacity of 128GB, its scalability limitations led to the lowest performance at the largest capacity of 1024GB, highlighting the drawbacks of conventional PIM mapping approaches at larger scales. In contrast, LoL-PIM demonstrated significantly better scalability. At 1024GB, it outperformed the baseline GPU-GDDR and PIM-only systems by 3.53$\times$ and 4.74$\times$, respectively, showcasing its superior efficiency. This improvement stems from two key factors: (1) pipeline parallelization, which sustains tensor parallelization and prevents PIM channel utilization drops at larger scales, and (2) larger batch sizes enabled by scaling, which reduce pipeline bubble overhead and enhance throughput  [^3] GPU-GDDR is evaluated to match the external bandwidth with PIM-only systems. A similar trend was observed with the Qwen1.5-72B model (Fig. 9(b)). The baseline PIM-only system again exhibited the lowest scalability and was unable to execute the 72B model on a 128GB system due to its parameter size. However, for capacities between 256GB and 1024GB, LoL-PIM consistently outperformed both the baseline GPU-GDDR and PIM-only systems by 8.54$\times$ and 2.65$\times$, respectively. These results underscore LoL-PIM’s capability to efficiently handle larger models with substantial memory and computational demands. Its sustained high throughput across increasing system capacities highlights its robustness and effectiveness in scaling large-scale workloads.

### v2

@@ 8.1. Evaluation Settings
Table 6. Validation of our Ramulator-based PIM simulator (cycle difference with the AiMX (Yongkee Kwon et al., 2024) vendor library).
@@
Table 7. Comparison of node specifications for GPU-HBM, GPU-GDDR, and PIM(=LoL-PIM).
@@
Baselines. We evaluate our system against three baselines, each representing the configuration of a single node, to ensure a fair and comprehensive comparison: (1) a DGX-A100 system, each GPU equipped with 80GB HBM3 and simulated using a DGX simulator (GPU-HBM), (2) AiMX cards (Yongkee Kwon et al., 2024), originally consisting of 2 modules, but expanded to 8 modules in our setup for fairness, with each module incorporating one Xilinx Virtex UltraScale+ VU9P FPGA(PIM Controller Hub) and eight 1GB AiM chips(2 channels per chip) utilizing GDDR6 and QSFP links for inter-card communication (PIM), and (3) a GPU-only system using GDDR6 for GPU evaluation with the matched external bandwidth of the PIM system (GPU-GDDR). These baselines were simulated using validated tools, including the DGX simulator (Jaewan Choi, 2024) and our extended version of Ramulator (Kim et al., 2016), which we validated to ensure accuracy (see Table 6). This setup enables a comprehensive comparison of LLM inference performance across architectures. Detailed specifications for the baselines are provided in Table 7.
@@ 8.2. Performance Comparison
Scalability. We evaluate LoL-PIM’s scalability across three long-context LLM benchmarks (Table 2), focusing on the Qwen1.5-7B model (Fig. 9(a)). The baseline PIM-only system showed the poorest scalability due to inefficient PIM utilization. While it outperformed the GPU-only system at a smaller capacity of 128GB, its scalability limitations led to the lowest performance at the largest capacity of 1024GB, highlighting the drawbacks of conventional PIM mapping approaches at larger scales. In contrast, LoL-PIM demonstrated significantly better scalability. At 1024GB, it outperformed the baseline GPU-GDDR and PIM-only systems by 3.53$\times$ and 4.74$\times$, respectively, showcasing its superior efficiency. This improvement stems from two key factors: (1) pipeline parallelization, which sustains tensor parallelization and prevents PIM channel utilization drops at larger scales, and (2) larger batch sizes enabled by scaling, which reduce pipeline bubble overhead and enhance throughput  [^3] GPU-GDDR is evaluated to match the external bandwidth with PIM-only systems. A similar trend was observed with the Qwen1.5-72B model (Fig. 10(a)). The baseline PIM-only system again exhibited the lowest scalability and was unable to execute the 72B model on a 128GB system due to its parameter size. However, for capacities between 256GB and 1024GB, LoL-PIM consistently outperformed both the baseline GPU-GDDR and PIM-only systems by 8.54$\times$ and 2.65$\times$, respectively. These results underscore LoL-PIM’s capability to efficiently handle larger models with substantial memory and computational demands. Its sustained high throughput across increasing system capacities highlights its robustness and effectiveness in scaling large-scale workloads.